OpenStudy (anonymous):

A sports company conducted a road test of a new model of a geared bike. The test rider cycled (3x - 2) miles on a flat road, (x2 - 5) miles uphill, and (2x + 7) miles downhill. Which simplified expression is equivalent to the total distance, in miles, for which the bike was tested? x2 - x x2 + 5x x2 − x + 14 x2 + 5x + 14 help me!!!!!!

OpenStudy (campbell_st):

well the total distance is the sum of the flat, uphill and downhill so \[D = 3x -2 + x^2 - 5 + 2x + 7\] just collect the like terms for the answer.

OpenStudy (anonymous):

that will be the 3x and 2x right

OpenStudy (anonymous):

do i add them????

OpenStudy (campbell_st):

yes as well as -2 -5 + 7

OpenStudy (campbell_st):

yep add 3x + 2x =

OpenStudy (anonymous):

3x+2x= 5

OpenStudy (campbell_st):

its 5x

OpenStudy (anonymous):

sorry 5x

OpenStudy (campbell_st):

yep.... now what about the numbers -2 -5 + 7 = and don't forget x^2 is sitting on its own...

OpenStudy (anonymous):

do i add them too??? and what do you mean x^2 is sitting on its own??????

OpenStudy (campbell_st):

well after 5x you have \[x^2 + 5x + 7 -2 - 5 \] you just need to simplify the last 3 numbers...

OpenStudy (anonymous):

um how do i do that lol

OpenStudy (campbell_st):

well what is 7 - 2= ?

OpenStudy (campbell_st):

then you need to subtract 5 from the answer (7 - 2) - 5 =?

OpenStudy (campbell_st):

then the question is all done... \[x^2 + 5x\]

OpenStudy (anonymous):

x+5=5 right

OpenStudy (anonymous):

7-2=5

OpenStudy (anonymous):

so x+5=5????

OpenStudy (secret-ninja):

Its B. :]